OpenStudy (anonymous):

Why is it false? The function of f(x)=2x-3/x-1 satisfies the MVT on the interval [0,2].

OpenStudy (anonymous):

f(2)-f(0)/2-1=-2 and is it because -2 isn't in the interval of [0,2]?

jimthompson5910 (jim_thompson5910):

it's (2x-3)/(x-1) right?

OpenStudy (anonymous):

yessir.

jimthompson5910 (jim_thompson5910):

ok, notice that when x = 0, f(x) is f(0) = (2*0-3)/(0-1) = (-3)/(-1) = 3 when x = 2, f(x) is f(2) = (2*2-3)/(2-1) = (1)/(1) = 1

jimthompson5910 (jim_thompson5910):

so we have these two points: (0,3) and (2,1) Find the slope of the line that is going through those two points for me please

OpenStudy (anonymous):

-1

jimthompson5910 (jim_thompson5910):

now the MVT states that if you have an interval [a,b] on some function f(x), then there exists some number c such that f ' (c) = ( f(b) - f(a) )/(b - a) Note: ( f(b) - f(a) )/(b - a) is the slope of the line you just calculated

OpenStudy (anonymous):

yea...

jimthompson5910 (jim_thompson5910):

so this means that if the MVT was true then there should be a value of c such that f ' (c) = -1

jimthompson5910 (jim_thompson5910):

can you find me such a value

OpenStudy (anonymous):

wait.. what? I'm confused now

jimthompson5910 (jim_thompson5910):

first off, what is f ' (x)

OpenStudy (anonymous):

1/(x-1)^2

jimthompson5910 (jim_thompson5910):

now solve 1/(c-1)^2 = -1 for c

jimthompson5910 (jim_thompson5910):

it might help to do this 1/(c-1)^2 = -1 1 = -1(c-1)^2 -1 = (c-1)^2 (c-1)^2 = -1

OpenStudy (anonymous):

c is a complex solution..

jimthompson5910 (jim_thompson5910):

so there are no real numbers c that satisfy f ' (c) = -1

OpenStudy (anonymous):

... can you sum all that up lol? Because I'm lost.

OpenStudy (anonymous):

So the slope is -1... But why does it not satisfy the thm?

jimthompson5910 (jim_thompson5910):

because there are no tangent lines that have a slope of -1 on [0,2]

jimthompson5910 (jim_thompson5910):

the secant slope is -1, so there should be a tangent line in [0,2] that also has a slope of -1 (if the MVT is correct)

OpenStudy (anonymous):

my gosh... I'm so lost! ;/ how do you know there are no tangent lines that have a slope of -1? is it because -1 isn't included in the interval?

jimthompson5910 (jim_thompson5910):

because f ' (c) = -1 has no real solutions

OpenStudy (anonymous):

oh so because f'(c)=-1 has no solutions it doesn't satisfy the thm.. But, if there were solutions it would satisfy the thm?

jimthompson5910 (jim_thompson5910):

yes, the big reason why this theorem fails is because of the jump discontinuity at x = 1

OpenStudy (anonymous):

@satellite73 what do you mean?

OpenStudy (anonymous):

you can see that the function does not exist at 1, so it sure as hell is not differentiable at 1

OpenStudy (anonymous):

hahaha^^!!